Wednesday, December 19, 2018

Night Log 2018-12-18

SA: Dani��l
SO: Thea
Others: Encarni, Richard, Timmy, Deneys (via phone), Hitesh (via phone)


The clouds cleared up at ~11pm, we observed 4 targets, one of which was a high priority. Humidity increased again at ~1 am, had to terminate the last observation. Humidity closed us for the rest of the evening.

Science data taken for:
2018-2-LSP-001 P0 RSS (Observing optical counterparts of transient objects)
2018-2-SCI-005 P3 HRS (Spectroscopy of TESS targets)
2017-1-MLT-004 P3 RSS (Monitoring the line emission in Southern Galactic Be/X-ray binaries)

No comments:

Post a Comment